OpenStudy (anonymous):

An internet search engine looks for a keyword in 9 databases, searching them in a random order. Only 5 of these databases contain the given keyword. Find the probability that it will be found in at least 2 of the first 4 searched databases. I'm kind of stuck setting up the problem.

OpenStudy (anonymous):

at least two out of four means 2 or 3 or 4 out of 4 so we can compute these probabilities and add or else compute the probability it is found in none or one of the first four and then subtract from one. your choice

OpenStudy (anonymous):

probability it is in all four is \[\frac{5}{9}\times \frac{4}{8}\times \frac{3}{7}\times \frac{2}{6}\] is that more or less clear?

OpenStudy (anonymous):

first and second and third and fourth, multiply the probabilities together

OpenStudy (anonymous):

or perhaps you might have been taught as follows all 4 means \[\frac{\dbinom{5}{4}}{\dbinom{9}{4}}\] you will get the same number

OpenStudy (anonymous):

then three out of four will be \[\frac{\dbinom{5}{3}\times \dbinom{4}{1}}{\dbinom{9}{4}}\]

OpenStudy (anonymous):

and finally 2 out of 4 will be \[\frac{\dbinom{5}{2}\times \dbinom{4}{2}}{\dbinom{9}{4}}\]

OpenStudy (anonymous):

do you know how to compute these numbers?

OpenStudy (anonymous):

I think I was over complicating the problem. I thought about computing the probability of at least 2 of the first 4 searches. Then I thought I needed to do more work since only 5 out of 9 databases contain the keyword. Now that you explained it, it makes sense. Thanks!

OpenStudy (anonymous):

yw but you do have to do a bit of computation

OpenStudy (anonymous):

Yes, I know how to perform the computations.

OpenStudy (anonymous):

ok good. have fun!

OpenStudy (anonymous):

Thanks, have a nice day!